# Task-Specific Directions: Definition, Exploration, and Utilization in Parameter Efficient Fine-Tuning

### v4

@@ 3.4 Challenges and Explorations: Using TSDs is Not as Easy as Expected
The results presented in Fig. 3 lead to two significant observations regarding the effectiveness of using $\Delta\mathbf{W}$ to identify TSDs:
@@ 5 How Much Can TSD Boost the Performance of LoRA?
To explore the performance gains of LoRA-Dash, LoRA-Init, and LoRA-TSD, both numerical and visual comparisons are employed. The numerical experiments focus on commonsense reasoning and natural language understanding tasks, assessing performance improvements of our methods in standard natural language processing metrics. Visual comparisons focuses on subject-driven generation task, involving using a text-to-image diffusion model to learn specific concepts (i.e., task-specific concept), providing a clear, qualitative and intuitive understanding of the impact of TSDs on performance. The detailed experimental settings are shown in Secs. 8.1 and 8.3-8.5, respectively.
@@ 5.1 Numerical Results
For comparisons with other methods, please refer to Tables. XI-XII.
@@ 5.2 Visual Results
5.3 Ablation Study of LoRA-Dash

### v5

@@ 3.4 Challenges and Explorations: Using TSDs is Not as Easy as Expected
The results are presented in Fig. 3, and we also present the statistic results at $r=32$ in Tables. XI-XII. It is obvious that the precision and recall of all layers remain consistently high for LoRA across different layers, training steps, and tasks. Since precision is computed over eight directional classes and recall over four, a mean precision of 75% corresponds to correctly identifying approximately six out of eight directions, while a recall of 75% indicates identifying three out of four. We also note that the reported standard deviations are relatively large, but this behavior is expected and statistically justified. For precision, misclassifying a single direction causes a drop of 12.5% in the metric, whereas for recall, each misclassification results in a 25% decrease. As the observed standard deviation for precision is close to this 12.5% unit drop—and the standard deviation for recall remains below the 25% threshold—we conclude that these variations fall within the normal statistical fluctuation induced by the discrete nature of directional classification. Moreover, we have also repeated the above experiment by changing the task to the natural language understanding task on the CoLA dataset. The results are shown in Tables. XIII-XIV. These results lead to two significant observations regarding the effectiveness of using $\Delta\mathbf{W}$ to identify TSDs:
@@ 5 How Much Can TSD Boost the Performance of LoRA?
To explore the performance gains of LoRA-Dash, LoRA-Init, and LoRA-TSD, both numerical and visual comparisons are employed. The numerical experiments focus on commonsense reasoning and natural language understanding tasks, assessing performance improvements of our methods in standard natural language processing metrics. Visual comparisons focuses on subject-driven generation task, involving using a text-to-image diffusion model to learn specific concepts (i.e., task-specific concept), providing a clear, qualitative and intuitive understanding of the impact of TSDs on performance.
5.1 Implementation Details
We compare our methods mainly with LoRA. The ranks of LoRA-Dash, LoRA-Init, LoRA-TSD, LoRA and other methods are varied among {2, 4, 8, 16, 32, 64}. All methods are implemented using the publicly available PyTorch [45] framework, and all experiments are conducted on NVIDIA A100 GPUs [^4] We have also test that most of the experiments can also be conducted on one consumer GPU resource such as NVIDIA RTX3090.. The hyper-parameter $t$ is set to 100, and $s=8$ for LoRA-TSD and LoRA-Dash. We fine-tune all the layers of different models in all the experiments, and the detailed experimental settings are shown in Secs. 8.2-8.4, respectively.
@@ 5.2 Numerical Results
For comparisons with other methods, please refer to Tables. XVIII-XIX.
@@ 5.3 Visual Results
5.4 Ablation Study and Sensitivity Analysis of LoRA-Dash
@@
Based on extensive empirical evaluations across multiple tasks, we suggest the following guidelines. First, the transition into the Dash phase should occur once training has reached a stable regime; in practice, this happens quickly, and we therefore recommend setting $t$ to a relatively small value, with a default of 100 steps, which is sufficient for convergence stabilization in most settings. Second, regarding the number of directions $s$, we use a default of 8. More complex tasks generally benefit from a larger number of directions, while simpler tasks tend to prefer fewer.